OpenStudy (anonymous):

Evaluate the derivative dy/dx at the (0,0) for the equation 2x -5x^3y^2+4y=0

hartnn (hartnn):

thats implicit derivative. do you know product and chain rules ?

OpenStudy (anonymous):

So I can just use the product rule?

hartnn (hartnn):

you will need both product and chain rule for the term 5x^3y^2

hartnn (hartnn):

let me give you example derivative of xy w.r.t x will be x (dy/dx) + y (dx/dx) = x dy/dx + y

OpenStudy (anonymous):

so is it 0= 2-5x^3* 2y -

OpenStudy (anonymous):

Sorry not that computer sent it when I wasnt done

OpenStudy (anonymous):

0= 2-5x^3 * 2y -15x^2 * y^2 + 4

hartnn (hartnn):

note : derivative of y^2 with respect to 'x' will require chain rule~ d/dx (y^2) = 2y dy/dx see if u get this ?

hartnn (hartnn):

similarly for 4y d/dx (4y) = 4 dy/dx (since we have 'y' but we are differentiating with respect to 'x', thats why we need chain rule)

OpenStudy (anonymous):

So wait it then becomes 2-5x^3 * 2 dy/dx - 15x^2 + 4 dy/dx

OpenStudy (anonymous):

If its not that, then Im completely lost and have no clue

hartnn (hartnn):

just one error, maybe typing, " 2-5x^3 * 2y dy/dx - 15x^2 + 4 dy/dx =0 "

hartnn (hartnn):

there should have been 2y dy/dx instead of 2 dy/dx because d/dx (y^2) = 2y dy/dx

hartnn (hartnn):

got it ?

OpenStudy (anonymous):

Oh okay yeah! I got it. So then do you leave it like that? or is the next step to make it all equal to dy/dt?

hartnn (hartnn):

we need to evaluate the derivative dy/dx AT (x,y) = (0,0) so, you just need to plug in x =0 and y = 0 in 2-5x^3 * 2y dy/dx - 15x^2 + 4 dy/dx = 0 and isolate dy/dx

OpenStudy (anonymous):

Sorry at last,there is a word 'near'.I didn't notice it before posting.

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (anonymous):

No problem...

OpenStudy (haseeb96):